Let xn = ( 2 + (−1)^n/ n+2 ). Using the algebra of limits and standard results on limits to establish/justify if the sequence {xn} n=1 to infinity converges or diverges; find the limit in case the sequence is convergent.



1
Expert's answer
2021-02-19T15:34:11-0500

"\lim \limits_{n \to \infty} \frac{2+(-1)^{n}}{n+2} = \lim \limits_{n \to \infty} \frac{\frac{2}{n} + \frac{(-1)^n}{n}}{\frac{n}{n}+ \frac{2}{n}}"

But

"\lim \limits_{n \to \infty} \frac{2}{n} = 0\\
\lim \limits_{n \to \infty} \frac{(-1)^n}{n} = 0"

Also

"\lim \limits_{n \to \infty} \frac{n}{n} = \lim \limits_{n \to \infty} 1 = 1"

"\lim \limits_{n \to \infty} \frac{2+(-1)^{n}}{n+2} = \frac{0+0}{1+0} = 0"

The sequence is convergent and the convergent point is zero.